Question 3

Needs Grading

GoM. had $50,000 in cash at year end 2011 and $25,000 in cash at year end 2012. The firm invested in Property and Plant and Equipment totaling $250,000. Cash Flow from financing totaled +$180,000.

What was the cash flow from operating activities?

If accruals increased by $30,000, receivables and inventories increased by $100,000, and depreciation and amortization totals $10,000, what was the firm’s net income?
Answer #1

Cash flow at year 2011 - Cash flow due to investing + Cash flow due to financing + cash flow from operation = Cash flow at year 2012

50000 - 250,000 + 180,000 + Cash flow from operations = 25000
Cash flow from operations = 25000+250000 -180000 -50000 = 45000

Net income + Increase in accruals - Increase in receivables and inventories + Depreciation = 45000
Net income + 30000 - 100000 + 10000 = 45000
Net income = 105000